1. 2x^2+10x+12
Factor completely
2.x^2+8x=-16

3.x^2-16=0
4. x^2-16x=-64
5. 2x^2-12x+10
6.x^2+8x+15=0
7. x^2-x-12=0

Answer by ankor@dixie-net.com(15646)

1. 2x^2+10x+12
factor out 2:
2(x^2 + 5x + 6)
2(x+3)(x+2)
:
2.x^2 + 8x = -16
x^2 + 8x + 16 = 0
(x+4)(x+4) = 0
:
3.x^2 - 16 = 0
Note the difference of squares
(x-4)(x+4)
:
4. x^2-16x=-64
x^2 - 16x + 64 = 0
(x-8)(x-8 = 0
:
5. 2x^2 - 12x + 10
Factor out 2
2(x^2 - 6x + 5)
2(x-5)(x-1)
:
6.x^2 + 8x + 15 = 0
(x+5)(x+3)
:
7. x^2 - x - 12 = 0
(x-4)(x+3)